Answer question NO.32) We showed a while back on a homework assignment that f(:) has a multiple root iff f(0) and f(:1) have a common factor. Th
Answer #1

f(x) = x²-3x+2 let #Given that Now ful= won²+x²-n-2n+2 = 0 n?-) ) + n(n-1) -2 (1-1) = 0 - (2-1) in ²+u-2) = 0 = (x-7) (n² + 2